OpenStudy (anonymous):

I need help with reversing the FOIL method. equation: 4x^3+3x^2+20x+15

OpenStudy (amistre64):

grouping ....

OpenStudy (amistre64):

4x^3+3x^2+20x+15 group it with ( )s (4x^3+3x^2) + (20x+15) now take out the commons for each set of ()

OpenStudy (anonymous):

dumb it down a little bit what do you mean by take out the commons ?

OpenStudy (amistre64):

what is the common factor of 4x^3 and 3x^2? what is the common factor of 20x and 15?

OpenStudy (anonymous):

greatest common factor lowest common factor what kind?

OpenStudy (amistre64):

thats a fair question ... we would want to take the greatest common factor in or to be able to reduce them to a simplest form

OpenStudy (amistre64):

*in order to ...

OpenStudy (anonymous):

I don't know how to do this im so lost.

OpenStudy (amistre64):

this would have been practiced a ton before ever getting to this type of question. I am assuming that you have been exposed to the simpler processes and I am simply trying to get you to apply what you would have already known.

OpenStudy (anonymous):

im just new to this Reverse FOIL thing

OpenStudy (amistre64):

what is the common factor of 4x^3 and 3x^2? 4xxx and 3xx have xx in common; so x^2 can be factored out (4x^3 + 3x^2) = x^2(4x+3) what is the common factor of 20x and 15? 4*5x and 3*5 have 5 in common, so we can factor that out (20x + 15) = 5(4x+3)

OpenStudy (amistre64):

FOIL is a memory device, and should never be taught in my opinion. If you know the distributive property, then you can work this all out

OpenStudy (amistre64):

the distributive property: a(x+y) = ax+ay if we have ax+ay, we can undistribute it by factoring out its common factor: ax+ay = a(x+y)

OpenStudy (amistre64):

so far we have worked it up as: \[4x^3+3x^2+20x+15\] \[(4x^3+3x^2)+(20x+15)\] \[x^2(4x+3)+5(4x+3)\] from here we should be able to see that both terms have (4x+3) in common; but if you want to see it better, let a = (4x+3) \[x^2a+5a\] in other words, lets "undistrbute" that a \[a(x^2+5)\] and since we defined a to be (4x+3) \[(4x+3)(x^2+5)\]

OpenStudy (anonymous):

I see how you got it. but lets practice another?

OpenStudy (amistre64):

one more then :)

OpenStudy (anonymous):

40n^3-20n^2+32n-16 lets take it slow step by step

OpenStudy (anonymous):

(40n^3-20n^2) (32n-16) ?

OpenStudy (amistre64):

dont run over that + sign v (40n^3-20n^2) + (32n-16) but yes, so far so good

OpenStudy (anonymous):

40nnn and 20nn simplifies to (40n+20)?

OpenStudy (amistre64):

go ahead and undistribute the nn, but there is still more they have in common; 40 = 20*2 40nnn and 20nn; 20nn in common 20n^2(2n-1) is better

OpenStudy (anonymous):

so now we gotta see what 32n and 16 have in common? which would be 8? which would give you 4n-2?

OpenStudy (anonymous):

no 2n-1

OpenStudy (amistre64):

2 and 4 still have a common factor so its not in simplest form yet 32n and 16 16*2n and 16n have 16n in common reducing to (2n-1) which is what we would expect to happen

OpenStudy (amistre64):

ugh, typed in that n by mistake

OpenStudy (amistre64):

edit**** 16*2n and 16n have 16 in common reducing to (2n-1) which is what we would expect to happen

OpenStudy (anonymous):

so now we have 20n^2(2n-1)(2n-1)?

OpenStudy (amistre64):

now lets see it in a big picture: \[40n^3-20n^2+32n-16\] \[(40n^3-20n^2)+(32n-16)\] \[20n^2(2n-1)+16(2n-1)\\~~~~~~~---~~~~~~~~~~---\\~~~~~~~~~~~~a~~~~~~~~~~~~~~~~~~~~a\]

OpenStudy (anonymous):

UGH I I don't know what to do on tis last step.

OpenStudy (amistre64):

since they have a common factor "a", undsitribute it

OpenStudy (amistre64):

\[20n^2(2n-1)+16(2n-1)\\~~~~~~~---~~~~~~~~~~---\\~~~~~~~~~~~~a~~~~~~~~~~~~~~~~~~~~a\] \[20n^2a+16a=a(20n^2+16)\]

OpenStudy (anonymous):

im lost on this last part with the "a"

OpenStudy (anonymous):

i actually have to go to lunch then switch classes ill be back at 12:15 eastern time ill come back. talk to you then. :(

OpenStudy (amistre64):

with a lot of practice on these, you can learn to see the common factor; im just trying to clean it up visually so that it is simpler to assess. the distributive property is defined as: a(x+y) = ax+ay, this is a fundamental property of algebra. we can "undistribute" ax+ay back into the form a(x+y) by factoring out the common "a"

OpenStudy (amistre64):

FOIL is a memory device that uses the distributive property: for example: (x+2) (5x-3) ; let a = (x+2) a(5x-3) = 5xa -3a we now have 2 more distributions to accomplish since a = (x+2) 5x(x+2) -3(x+2) 5x(x) + 5x(2) -3(x) -3(2) 5x^2 + 10x -3x - 6 5x^2 + (10-3)x - 6 5x^2 + 7x - 6

OpenStudy (anonymous):

Here is a link: http://www.youtube.com/watch?v=162ub1GIuqA. It's the same explanation as amistre64 but it might help you to see someone explaining it as a video. Also you might want to look up factoring using polynomial long division. Sometimes it's easier to factor a third degree polynomial that way.